A company has two products: A and B. It uses activity-based costing and has prepared the following analysis showing budgeted cost and activity for each of its three activity cost pools:

A company has two products: A and B. It uses activity-based costing and has prepared the following analysis showing budgeted cost and activity for each of its three activity cost pools:

Budgeted Activity
Activity Cost
Pool
Budgeted
Cost
Product A Product B
Activity 1 $ 106,000 4,900 4,700
Activity 2 $ 81,000 6,400 7,400
Activity 3 $ 131,000 4,400 7,150


Annual production and sales level of Product A is 36,200 units, and the annual production and sales level of Product B is 71,450 units. What is the approximate overhead cost per unit of Product A under activity-based costing?

Multiple Choice

  • $1.41

  • $3.91

  • $2.47

  • $10.26

  • $11.04


Annual production and sales level of Product A is 36,200 units, and the annual production and sales level of Product B is 71,450 units. What is the approximate overhead cost per unit of Product A under activity-based costing?
Answer #1

B. $3.91

Activity 1 allocated to Product A line = $106,000 × 4,900/9,600 = $54,104

Activity 2 allocated to Product A line: $81,000 × 6,400/13,800 = $37,565

Activity 3 allocated to Product A line: $131,000 × 4,400/11,550 = $49,905

Total overhead allocated to Product A = $141,574

Overhead per unit of Product A: $141,574/36,200 = $3.91